ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 245, 2021

Araújo, D. J., E. V. Teixeira and J. M. Urbano On a two-phase free boundary problem ruled by the infinity Laplacian

773

Avvakumov, S., I. Mabillard, A. B. Skopenkov and U. Wagner Eliminating higher-multiplicity intersections. III. Codimension 2

501

Balitskiy, A. A multi-plank generalization of the Bang and Kadets inequalities

209

Bergelson, V. and Ferré Moragues, A. An ergodic correspondence principle, invariant means and applications

921

Carvalho, M. and S. A. Pérez Periodic points and measures for a class of skew-products

455

Chirvasitu, A., F. Ladisch and P. Soberón Finite groups as prescribed polytopal symmetries

75

Ciobanu, L. and M. Elder The complexity of solution sets to equations in hyperbolic groups

869

Ciosmak, K. J. Continuity of extensions of Lipschitz maps  567

D’Angeli, D., E. Rodaro and J. P. Wächter Corrigendum to “Automaton semigroups and groups: On the undecidability of problems related to freeness and finiteness”

535

Dolinka, I. Free idempotent generated semigroups: The word problem and structure via gain graphs

347

Džamonja, M. and J. Väänänen Chain logic and Shelah’s infinitary logic

93

Elder, M. See L. Ciobanu

Fernandes, G., M. Moreno and A. Rinot Fake reflection  295

A. Ferré Moragues See V. Bergelson Figueiredo, G. and M. Montenegro FitzHugh–Nagumo system

with zero mass and critical growth

711

Gishboliner, L. and A. Shapira Testing graphs against an unknown distribution

787

Hartz, M. and M. Lupini Dilation theory in finite dimensions and matrix convexity

39

Kaya, B. and I. Sağlam On the closed Ramsey numbers R^{cl}(ω+n, 3)  963

Kerman, E. and Y. Liang Higher index symplectic capacities do not satisfy the symplectic Brunn–Minkowski inequality 27 Kleinbock, D., N. Moshchevitin and B. Weiss Singular vectors

on manifolds and fractals

589

Ladisch, F. See A. Chirvasitu Liang, Y. See E. Kerman Lupini, M. See M. Hartz

Ma, J. and T. Yang Non-repeated cycle lengths and Sidon sequences  639

Mabillard, I. See S. Avvakumov Minamoto, H. Resolutions and homological dimensions of DG-

modules

409

Minasyan, A. Some examples of invariably generated groups  231

Montenegro, M. See G. Figueiredo Moreno, M. See G. Fernandes Moshchevitin, N. See D. Kleinbock Pérez, S. A. See M. Carvalho Pham, L. and F. Thilmany Arithmetic groups and the Lehmer con-

jecture

839

Phillips, A. Impossible configurations for geodesics on negativelycurved surfaces

165

Pimentel, E. A. and M. S. Santos Improved regularity for the porous medium equation along zero level sets

135

Qiu, Y. and K. Wang Rigidity of determinantal point processes on the unit disc with sub-Bergman kernels

389

Ren, Y. Quadratic torsion subgroups of modular Jacobian varieties  675

Revin, D. O. See N. Yang Rinot, A. See G. Fernandes Rodaro, E. See D. D’Angeli Sağlam, I. See B. Kaya Santos, M. S. See E. A. Pimentel Shapira, A. See L. Gishboliner

Shen, W. and Y. Wang Primitive tuning via quasiconformal surgery  259

Shestakov, I. and M. Zaicev Codimension growth of simple Jordan superalgebras

615

Skopenkov, A. B. See S. Avvakumov

Skryabin, S. Flatness over PI coideal subalgebras  735

Soberón, P. See A. Chirvasitu Teixeira, E. V. See D. J. Araújo Thilmany, F. See L. Pham Urbano, J. M. See D. J. Araújo Väänänen, J. See M. Džamonja Vdovin, E. P. See N. Yang Wächter, J. P. See D. D’Angeli Wagner, U. See S. Avvakumov Wang, K. See Y. Qiu

Wang, Q. and Z. Wang Notes on the ℓ^{p}-Toeplitz algebra on ℓ^{p}(ℕ)  153

Wang, Y. See W. Shen Wang, Z. See Q. Wang Weiss, B. See D. Kleinbock

Xie, J. and Z. Xu Subset selection for matrices with fixed blocks  1

Xu, Z. See J. Xie Xue, L. A Proof of Grünbaum’s Lower Bound Conjecture for general

polytopes

991

Yakir, O. and O. Zeitouni The minimum modulus of Gaussian trigonometric polynomials

543

Yang, N., D. O. Revin and E. P. Vdovin Baer–Suzuki theorem for the π-radical

173

Yang, T. See J. Ma Zaicev, M. See I. Shestakov Zeitouni, O. See O. Yakir